OpenStudy (anonymous):

What is the 19th term of the arithmetic sequence 15, 23, 31, ...?

OpenStudy (anonymous):

y=mx+b 23=2m+b 15=m+b When you solve for m and b m=8 b=7 Now you have your equation y=8x+7 You can plug in your input, which is 19 y=8*19+7 y=159 the answer will be 159

OpenStudy (anonymous):

or just add 8 to the number...

OpenStudy (anonymous):

It would kind of take a long time to add 8 to each number in the sequence up to the 19th term